# A Comprehensive Survey on Vector Database: Storage and Retrieval Technique, Challenge

### v2

@@ II-C Caching
Least Frequently Used (LFU). The LFU algorithm is a frequency-based cache eviction mechanism that determines removal priority by continuously tracking the access count of each vector data item. The LFU algorithm maintains a frequency table and evicts the least frequently accessed items when the cache is full. A typical implementation requires maintaining an access counter for each cached item, often using a min-heap data structure to efficiently identify the lowest-frequency items with a time complexity of O(log n)[45, 46]. However, the LFU algorithm can be susceptible to cache pollution, as items frequently accessed in the past may linger in the cache even when they are no longer needed. Thus, LFU is suitable for applications with stable and long-lived spot data patterns (e.g., popular product recommendations, high-frequency user profile queries), while its effectiveness diminishes in environments with rapidly evolving access distributions.
Partitioned Cache. Partitioned caching is a common approach in VDBs, wherein vector data are divided into multiple partitions based on specific criteria [47], such as geographic location, category, or access frequency. Each partition can be allocated a distinct cache size and may employ different eviction policies according to the respective demand and usage patterns. This enables the cache to retain the most relevant vector data for each partition, thereby improving resource utilization and overall cache effectiveness. For example, Esri, a leading geographic information system (GIS) company, leverages the partitioned cache to efficiently store vector data and support high-performance map rendering.
II-D Replication
@@
Replication is a technique that creates multiple copies of the vector data and stores them on different nodes or clusters. Replication can improve the availability, durability, and performance of VDB. This section discusses three common replication methods: leader-follower replication, multi-leader replication, and leaderless replication.
@@
Leader-Follower Replication. Leader-Follower replication designates one node as the leader and the others as the followers, and allows only the leader to accept write requests and propagate them to the followers[48]. Leader-follower replication can ensure strong consistency and simplify the conflict resolution of VDB. However, it may also introduce availability issues and require failover mechanisms to handle leader failures.
@@
Multi-Leader Replication. Multi-Leader replication extends the traditional leader-follower model by designating multiple nodes as leaders, each capable of independently accepting and processing write requests[49, 50]. In this architecture, all leader nodes can concurrently handle write operations and asynchronously propagate changes to other nodes in the system.
@@
Leaderless Replication. Leaderless replication does not distinguish between leader and follower nodes, and allows any node to accept write and read requests [51, 52]. Leaderless replication can avoid single points of failure and improve the scalability and reliability of VDB. However, it may also introduce consistency issues and require coordination mechanisms to resolve conflicts.
III Search
@@
Observing the current division for NNS and ANNS algorithms, the boundary is precisely their design principle, such as how they organize, index, or hash the dataset, how they search or traverse the data structure, and how they measure or estimate the distance between points. NNS algorithms tend to use more exact or deterministic methods, such as partitioning the space into regions by splitting along one dimension (k-d tree) or enclosing groups of points in hyperspheres (ball tree), and visiting only the regions that may contain the nearest neighbor based on some distance bounds or criteria. ANNS algorithms tend to use more probabilistic or heuristic methods, such as mapping similar points to the same or nearby buckets with high probability (locality-sensitive hashing), visiting the regions in order of their distance to the query point and stopping after a fixed number of regions or points (best bin first), or following the edges that lead to closer points in a graph with different levels of coarseness (hierarchical navigable small world).
In fact, a data structure or algorithm that supports NNS can also be applied to ANNS, and for ease of categorization, such methods are included under the section on NNS. And in recent years, several new algorithms for high-dimensional vector NNS have emerged [53, 54, 55, 56, 13]. Although these algorithms have not yet been widely adopted by VDBs, they hold significant potential for future applications.
III-A Nearest Neighbor Search
III-A1 Brute Force Approach
A brute force algorithm for the NNS problem scans all points in the dataset, computing their distances to the query point and tracking the closest one. This algorithm guarantees to find the true nearest neighbor for any query point, but it has a high computational cost. The time complexity of a brute force algorithm for NNS problem is $O(n)$, where n is the size of the dataset. The space complexity is $O(1)$, since no extra space is needed.
III-A2 Tree-Based Approach
@@
KD-Tree[57] . It is a technique for organizing points in a k-dimensional space, where k is usually a very big number. It works by building a binary tree in which every node is a k-dimensional point. Every non-leaf node in the tree acts as a splitting hyperplane that divides the space into two parts, known as half-spaces. The splitting hyperplane is perpendicular to the chosen axis, which is associated with one of the k dimensions. The splitting value is usually the median or the mean of the points along that dimension.
The algorithm maintains a priority queue of nodes to visit, sorted by their distance to the query point. At each step, the algorithm pops the node with the smallest distance from the queue, and checks if it is a leaf node or an internal node. If it is a leaf node, the algorithm compares the distance between the query point and the data point stored in the node, and updates the current best distance and nearest neighbor if necessary. If it is an internal node, the algorithm pushes its left and right children to the queue, with their distances computed as follows:
[…]
@@
The advantage of KD-tree is that it is conceptually simpler and often easier to implement than some of the other tree structures. The performance of KD-tree depends on several factors, such as the dimensionality of the space, the number of points, and the distribution of the points. These factors affect the trade-off between accuracy and efficiency, as well as the complexity and scalability of the algorithm. There are also some challenges and extensions of KD-tree, such as dealing with the curse of dimensionality when the dimensionality is high, introducing randomness in the splitting process to improve robustness, or using multiple trees to increase recall. This is a variation of KD-tree named randomized KD-tree, that introduces some randomness in the splitting process, which can improve the performance of KD-tree by reducing its sensitivity to noise and outliers [58].
Ball-Tree [59], [60], [61]. It is a technique for finding the nearest neighbors of a given vector in a large collection of vectors. It works by building a ball-tree, which is a binary tree that partitions the data points into balls, i.e. hyperspheres that contain a subset of the points. Each node of the ball-tree defines the smallest ball that contains all the points in its subtree. The algorithm then searches for the closest ball to the query point, and then searches within the closest ball to find the closest point to the query point.
To query for the nearest neighbor of a given point, the ball tree algorithm uses a priority queue to store the nodes to be visited, sorted by their distance to the query point. The algorithm starts from the root node and pushes its two children to the queue. Then, it pops the node with the smallest distance from the queue and checks if it is a leaf node or an internal node. If it is a leaf node, it computes the distance between the query point and each data point in the node, and updates the current best distance and nearest neighbor if necessary. If it is an internal node, it pushes its two children to the queue, with their distances computed as follows:
$\begin{array}[]{l}d_{L}(q,N)=\max(0,{N.value}-\|q-{N.center}\|)\\ d_{R}(q,N)=\max(0,\|q-{N.center}\|-N.value)\end{array}$ (3)
@@
The advantage of ball-tree is that it can perform well in high-dimensional spaces, as it can avoid the curse of dimensionality that affects other methods such as KD-tree. The performance of ball-tree depends on several factors, such as the dimensionality of the data, the number of balls per node, and the distance approximation method used. These factors affect the trade-off between accuracy and efficiency, as well as the complexity and scalability of the algorithm. There are also some challenges and extensions of ball-tree search, such as dealing with noisy and outlier data, choosing a good splitting dimension and value for each node, or using multiple trees to increase recall.
R-Tree [62]. It is a technique for finding the nearest neighbors of a given vector in a large collection of vectors. It works by building an R-tree, which is a tree data structure that partitions the data points into rectangles, i.e. hyperrectangles that contain a subset of the points. Each node of the R-tree defines the smallest rectangle that contains all the points in its subtree. The algorithm then searches for the closest rectangle to the query point, and then searches within the closest rectangle to find the closest point to the query point.
@@
$MBR(P)=\left[\min_{p\in P}p_{x},\max_{p\in P}p_{x}\right]\times\left[\min_{p\in P}p_{y},\max_{p\in P}p_{y}\right]$ (4)
where $p_{x}$ and $p_{y}$ are the $\mathrm{x}$ and $\mathrm{y}$ coordinates of point $p$ , and $\times$ denotes the Cartesian product.
The R-tree algorithm also uses two metrics to measure the quality of a node split: area and overlap. The area of a node is the area of its MBR, and the overlap of two nodes is the area of the intersection of their MBRs. The formula for computing the area of a node N is:
$\operatorname{area}(N)=\left(N.x_{\max}-N.x_{\min}\right)\times\left(N\cdot y_{\max}-N\cdot y_{\min}\right)$ (5)
where $N.x_{\min}$, $N.x_{\max}$, $N.y_{\min}$, and $N.y_{\max}$ are the coordinates of the MBR of node $N$.
The advantage of R-tree is that it can support spatial queries, such as range queries or nearest neighbor queries, on data points that represent geographical coordinates, rectangles, polygons, or other spatial objects. R-tree search performance depends on roughly the same factors as B-tree, and also faces similar challenges as B-tree.
M-Tree [63]. It is a technique for finding the nearest neighbors of a given vector in a large collection of vectors. It works by building an M-tree, which is a tree data structure that partitions the data points into balls, i.e. hyperspheres that contain a subset of the points. Each node of the M-tree defines the smallest ball that contains all the points in its subtree. The algorithm then searches for the closest ball to the query point, and then searches within the closest ball to find the closest point to the query point.
The M-tree algorithm uses the concept of covering radius to represent the spatial objects in the tree. The covering radius of a node is the maximum distance from the node’s routing object to any of its children objects. The formula for computing the covering radius of a node $N$ is:
$r(N)=\max_{C\in N.child}d(N.object,C.object)$ (6)
@@
$\operatorname{area}(N)=\sum_{C\in\text{ N.child }}\pi r(C)^{2}$ (7)
where $\pi$ is the mathematical constant, and $r(C)$ is the covering radius of child node $C$.
The advantage of M-tree is that it can support dynamic operations, such as inserting or deleting data points, by updating the tree structure accordingly. M-tree search performance depends on roughly the same factors as B-tree, and also faces similar challenges as B-tree.
III-B Approximate Nearest Neighbor Search
III-B1 Hash-Based Approach
The core idea of the hash-based approach is to reduce search complexity by mapping high-dimensional data to lower-dimensional hash codes with carefully designed hash functions, while preserving similarity between data points. As shown in Figure 3, each high-dimensional vector is transformed into a low-dimensional hash code. Similar points are mapped to the same or neighboring codes, so the search only needs to examine a small subset of codes, greatly improving efficiency. Based on this idea, four representative methods will be introduced: locality-sensitive hashing, spectral hashing, Spherical Hashing, and deep hashing. The idea is to reduce the memory footprint and the search time by comparing the binary codes instead of the original vectors [64].
@@ III-B1 Hash-Based Approach
Local-Sensitive Hashing [65, 66]. It is a technique for finding the approximate nearest neighbors of a given vector in a large collection of vectors. It works by using a hash function to transform the high-dimensional vectors into compact binary codes, and then using a hash table to store and retrieve the codes based on their similarity or distance. In LSH, hash functions are designed to preserve the locality of vectors. Unlike traditional hash functions, LSH increases the probability that similar items are mapped to the same code, thus increasing collisions among similar vectors. A trace of algorithm description and implementation for locally sensitive hashing can be seen on the home page [67].
The LSH algorithm works by using a family of hash functions that use random projections or other techniques which are locality sensitive, meaning that similar vectors are more likely to have the same or similar codes than dissimilar vectors [68], which satisfy the following property:
@@
$f(d(p,q))=1-\frac{d(p,q)}{\pi w}$ (10)
where $d(p,q)$ is the Euclidean distance between $p$ and $q$. The advantage of LSH is that it can reduce the memory footprint and the search time by comparing the binary codes instead of the original vectors, also adapt to dynamic data sets, by inserting or deleting codes from the hash table without affecting the existing codes [69].
The performance of LSH depends on several factors, such as the dimensionality of the data, the number of hash functions, the number of bits per code, and the desired accuracy and recall. These factors affect the trade-off between accuracy and efficiency, as well as the complexity and scalability of the algorithm. There are also some challenges and extensions of LSH, such as dealing with noisy and outlier data, choosing a good hash function family, or using multiple hash tables to increase recall. It is improved by [70], [71], [72].
Spectral Hashing [73]. It is a technique for finding the approximate nearest neighbors of a given vector in a large collection of vectors. It works by using spectral graph theory to generate hash functions that minimize the quantization error and maximize the variance of the binary codes. Spectral hashing can perform well when the data points lie on a low-dimensional manifold embedded in a high-dimensional space.
@@
$\min_{y_{1},\ldots,y_{n}}\sum_{i=1}^{n}Var\left(y_{i}\right)-\lambda I\left(y_{1},\ldots,y_{n}\right)$ (11)
@@
Spherical Hashing. Spherical hashing is a binary encoding technique based on hyperspheres, designed for efficient ANNS. Unlike traditional hyperplane-based methods, it partitions the data space using hyperspheres, which define tighter and more compact regions through their centers and radii. Each spherical hashing function, as described by Heo [74], is characterized by $(p_{k}\in\mathbb{R}^{D})$ and a distance threshold $t_{k}\in\mathbb{R}^{+}$, as detailed below:
$h_{k}(x)=\begin{cases}-1&\text{ when }d\left(p_{k},x\right)>t_{k}\\ +1&\text{ when }d\left(p_{k},x\right)\leq t_{k}\end{cases}$ (12)
@@
Deep Hashing [75, 76]. It is a technique for finding the approximate nearest neighbors of a given vector in a large collection of vectors. It works by using a deep neural network to learn hash functions that transform high-dimensional vectors into compact binary codes, and then using a hash table to store and retrieve the codes based on their similarity or distance[77]. The hash functions are designed to preserve the semantic information of the vectors, which means that similar vectors are more likely to have the same or similar codes than dissimilar vectors [78].
@@
$\min_{W,B}\sum_{i=1}^{N}\left\|x_{i}-Wb_{i}\right\|_{2}^{2}+\lambda\left\|b_{i}-\operatorname{sgn}\left(b_{i}\right)\right\|_{2}^{2}$ (14)
@@
The advantage of deep hashing is that it can leverage the representation learning ability of neural networks to generate more discriminative and robust codes for complex data, such as images, texts, or audios. The performance of deep hashing depends on several factors, such as the architecture of the neural network, the loss function used to train the network, and the number of bits per code. These factors affect the trade-off between accuracy and efficiency, as well as the complexity and scalability of the algorithm. There are also some challenges and extensions of deep hashing, such as dealing with noisy and outlier data, choosing a good initialization for the network, or using multiple hash functions to increase recall.
III-B2 Tree-Based Approach
The main idea of the tree-based approach is to build hierarchical or recursively partitioned data structures, such as trees, to break high-dimensional datasets into smaller subsets. This method improves query efficiency by reducing the number of points that need to be searched. Along this line, three tree-based methods will be presented: approximate nearest neighbors oh yeah, best bin first, and k-means tree. The idea is to reduce the search space by following the branches of the tree that are most likely to contain the nearest neighbors of the query point.
Approximate Nearest Neighbors Oh Yeah [79]. It is a technique which can perform fast and accurate similarity search and retrieval of high-dimensional vectors. It works by building a forest of binary trees, where each tree splits the vector space into two regions based on a random hyperplane. Each vector is then assigned to a leaf node in each tree based on which side of the hyperplane it falls on. To query a vector, Annoy traverses each tree from the root to the leaf node that contains the vector, and collects all the vectors in the same leaf nodes as candidates. Then, it computes the exact distance or similarity between the query vector and each candidate, and returns the top $k$ nearest neighbors.
The formula for finding the median hyperplane between two points $p$ and $q$ is:

### v3

@@ 2.3 Caching
Least Frequently Used (LFU). The LFU algorithm is a frequency-based cache eviction mechanism that determines removal priority by continuously tracking the access count of each vector data item. The LFU algorithm maintains a frequency table and evicts the least frequently accessed items when the cache is full. A typical implementation requires maintaining an access counter for each cached item, often using a min-heap data structure to efficiently identify the lowest-frequency items with a time complexity of O(log n) [lee1999existence, podlipnig2003survey]. However, the LFU algorithm can be susceptible to cache pollution, as items frequently accessed in the past may linger in the cache even when they are no longer needed. Thus, LFU is suitable for applications with stable and long-lived spot data patterns (e.g., popular product recommendations, high-frequency user profile queries), while its effectiveness diminishes in environments with rapidly evolving access distributions.
Partitioned Cache. Partitioned caching is a common approach in VDBs, wherein vector data are divided into multiple partitions based on specific criteria [mittal2017survey], such as geographic location, category, or access frequency. Each partition can have its own size and eviction policy. This method offers greater flexibility and better aligns caching behavior with application needs. For example, geographic information systems (GIS) often partition vector data by region to support fast and localized map rendering. When configured well, partitioned caches can improve hit rates while preventing a few high-frequency.
2.4 Replication
@@
Replication means making several copies of vector data and placing them on different nodes or clusters. This helps boost the availability, reliability, and overall performance of a VDB. In this section, we look at three widely used replication strategies: leader-follower, multi-leader, and leaderless replication.
@@
Leader-Follower Replication. Leader-Follower replication designates one node as the leader and the others as the followers and allows only the leader to accept write requests and propagate them to the followers [ongaro2014search]. Leader-follower replication can ensure strong consistency and simplify the conflict resolution of VDB. However, it may also introduce availability issues and require failover mechanisms to handle leader failures.
@@
Multi-Leader Replication. Multi-Leader replication extends the traditional leader-follower model by designating multiple nodes as leaders, each capable of independently accepting and processing write requests [garcia1985assign, garmany2003oracle]. In this architecture, all leader nodes can concurrently handle write operations and asynchronously propagate changes to other nodes in the system.
@@
Leaderless Replication. Leaderless replication does not distinguish between leader and follower nodes and allows any node to accept write and read requests [decandia2007dynamo, bailis2014quantifying]. Leaderless replication can avoid single points of failure and improve the scalability and reliability of VDB. However, it may also introduce consistency issues and require coordination mechanisms to resolve conflicts.
3 Search
@@
NNS algorithms tend to use more exact or deterministic methods, such as partitioning the space into regions by splitting along one dimension (k-d tree) or enclosing groups of points in hyperspheres (ball tree), and visiting only the regions that may contain the nearest neighbor based on some distance bounds or criteria. ANNS algorithms tend to use more probabilistic or heuristic methods, such as mapping similar points to the same or nearby buckets with high probability (locality-sensitive hashing), visiting the regions in order of their distance to the query point and stopping after a fixed number of regions or points (best bin first), or following the edges that lead to closer points in a graph with different levels of coarseness (hierarchical navigable small world).
In fact, a data structure or algorithm that supports NNS can also be applied to ANNS, and for ease of categorization, such methods are included under the section on NNS. And in recent years, several new algorithms for high-dimensional vector NNS have emerged [gao2023high, manohar2024parlayann, fu2021high, wang2023dumpy, zhao2023towards]. Although these algorithms have not yet been widely adopted by VDBs, they hold significant potential for future applications.
3.1 Nearest Neighbor Search
3.1.1 Brute Force Approach
A brute force algorithm for the NNS problem scans all points in the dataset, computing their distances to the query point and tracking the closest one. This algorithm guarantees to find the true nearest neighbor for any query point, but it has a high computational cost. The time complexity of a brute force algorithm for the NNS problem is $O(n)$, where n is the size of the dataset. The space complexity is $O(1)$, since no extra space is needed.
3.1.2 Tree-Based Approach
@@
KD-Tree [bentley1975multidimensional]. It is a technique for organizing points in a k-dimensional space, where k is usually a very big number. It works by building a binary tree in which every node is a k-dimensional point. Every non-leaf node in the tree acts as a splitting hyperplane that divides the space into two parts, known as half-spaces. The splitting hyperplane is perpendicular to the chosen axis, which is associated with one of the k dimensions. The splitting value is usually the median or the mean of the points along that dimension.
The algorithm maintains a priority queue of nodes to visit, sorted by their distance to the query point. At each step, the algorithm pops the node with the smallest distance from the queue and checks if it is a leaf node or an internal node. If it is a leaf node, the algorithm compares the distance between the query point and the data point stored in the node and updates the current best distance and nearest neighbor if necessary. If it is an internal node, the algorithm pushes its left and right children to the queue, with their distances computed as follows:
$\vskip-8.53581pt\small d_{L}(q,N)=\begin{cases}0&\text{if }q_{\mathit{N.axis}}\leq N.\mathit{value}\\ (q_{\mathit{N.axis}}-N.\mathit{value})^{2}&\text{if }q_{\mathit{N.axis}}>N.\mathit{value}\end{cases}$ (1)
$\vskip-8.53581pt\small d_{R}(q,N)=\begin{cases}0&\text{if }q_{\mathit{N.axis}}\geq N.\mathit{value}\\ (N.\mathit{value}-q_{\mathit{N.axis}})^{2}&\text{if }q_{\mathit{N.axis}}<N.\mathit{value}\end{cases}$ (2)
@@
The advantage of KD-tree is that it is conceptually simpler and often easier to implement than some of the other tree structures. The performance of KD-tree depends on several factors, such as the dimensionality of the space, the number of points, and the distribution of the points. There are also some challenges and extensions of KD-tree, such as dealing with the curse of dimensionality when the dimensionality is high, introducing randomness in the splitting process to improve robustness, or using multiple trees to increase recall. This is a variation of KD-tree named randomized KD-tree that introduces some randomness in the splitting process, which can improve the performance of KD-tree by reducing its sensitivity to noise and outliers [ghojogh2018treebased].
Ball-Tree [omohundro1989five, liu2006new, dolatshah_ball-tree_2015]. It is a technique for finding the nearest neighbors of a given vector in a large collection of vectors. It works by building a ball-tree, which is a binary tree that partitions the data points into balls, i.e., hyperspheres that contain a subset of the points. Each node of the ball-tree defines the smallest ball that contains all the points in its subtree. The algorithm then searches for the closest ball to the query point and then searches within the closest ball to find the closest point to the query point.
To query for the nearest neighbor of a given point, the ball tree algorithm uses a priority queue to store the nodes to be visited, sorted by their distance to the query point. The algorithm starts from the root node and pushes its two children to the queue. Then, it pops the node with the smallest distance from the queue and checks if it is a leaf node or an internal node. If it is a leaf node, it computes the distance between the query point and each data point in the node and updates the current best distance and nearest neighbor if necessary. If it is an internal node, it pushes its two children to the queue, with their distances computed as follows:
$\vskip-2.84526pt\small\begin{array}[]{l}d_{L}(q,N)=\max(0,{N.value}-\|q-{N.center}\|)\\ d_{R}(q,N)=\max(0,\|q-{N.center}\|-N.value)\end{array}$ (3)
@@
The advantage of ball-tree is that it can perform well in high-dimensional spaces, as it can avoid the curse of dimensionality that affects other methods , such as KD-tree. The performance of ball-tree depends on several factors, such as the dimensionality of the data, the number of balls per node, and the distance approximation method used. There are also some challenges and extensions of ball-tree search, such as dealing with noisy and outlier data, choosing a good splitting dimension and value for each node, or using multiple trees to increase recall.
R-Tree [guttman1984r]. It is a technique for finding the nearest neighbors of a given vector in a large collection of vectors. It works by building an R-tree, which is a tree data structure that partitions the data points into rectangles, i.e., hyperrectangles that contain a subset of the points. Each node of the R-tree defines the smallest rectangle that contains all the points in its subtree. The algorithm then searches for the closest rectangle to the query point and then searches within the closest rectangle to find the closest point to the query point.
@@
$\vskip-5.69054pt\small MBR(P)=\left[\min_{p\in P}p_{x},\max_{p\in P}p_{x}\right]\times\left[\min_{p\in P}p_{y},\max_{p\in P}p_{y}\right]$ (4)
where $p_{x}$ and $p_{y}$ are the $\mathrm{x}$ and $\mathrm{y}$ coordinates of point $p$ , and $\times$ denotes the Cartesian product. The R-tree algorithm also uses two metrics to measure the quality of a node split: area and overlap. The area of a node is the area of its MBR, and the overlap of two nodes is the area of the intersection of their MBRs. The formula for computing the area of a node N is:
$\vskip-2.84526pt\small\operatorname{area}(N)=\left(N.x_{\max}-N.x_{\min}\right)\times\left(N\cdot y_{\max}-N\cdot y_{\min}\right)$ (5)
where $N.x_{\min}$, $N.x_{\max}$, $N.y_{\min}$, and $N.y_{\max}$ are the coordinates of the MBR of node $N$. The advantage of R-tree is that it can support spatial queries, such as range queries or nearest neighbor queries, on data points that represent geographical coordinates, rectangles, polygons, or other spatial objects. R-tree search performance depends on roughly the same factors as B-tree and also faces similar challenges as B-tree.
M-Tree [ciaccia1997m]. It is a technique for finding the nearest neighbors of a given vector in a large collection of vectors. It works by building an M-tree, which is a tree data structure that partitions the data points into balls, i.e., hyperspheres that contain a subset of the points. Each node of the M-tree defines the smallest ball that contains all the points in its subtree. The algorithm then searches for the closest ball to the query point and then searches within the closest ball to find the closest point to the query point.
The M-tree algorithm uses the concept of covering radius to represent the spatial objects in the tree. The covering radius of a node is the maximum distance from the node’s routing object to any of its child objects. The formula for computing the covering radius of a node $N$ is:
$\vskip-2.84526ptr(N)=\max_{C\in N.child}d(N.object,C.object)$ (6)
@@
$\vskip-2.84526pt\operatorname{area}(N)=\sum_{C\in\text{ N.child }}\pi r(C)^{2}$ (7)
where $\pi$ is the mathematical constant, and $r(C)$ is the covering radius of child node $C$. The advantage of M-tree is that it can support dynamic operations, such as inserting or deleting data points, by updating the tree structure accordingly. M-tree search performance depends on roughly the same factors as B-tree, and also faces similar challenges as B-tree.
3.1.3 Summary and Trade-offs of NNS Methods
Exact nearest neighbor search guarantees returning the true nearest neighbor but faces performance challenges as data size and dimensionality grow. Brute-force search is simple and accurate but its linear complexity limits it to small datasets.
Tree-based methods (KD-Tree, Ball-Tree, R-Tree, M-Tree) reduce search space through spatial partitioning. KD-Tree works well in low dimensions but suffers from the curse of dimensionality. Ball-Tree handles high dimensions better at higher construction cost. R-Tree supports rectangle queries for spatial data but may require multiple search paths due to node overlap. M-Tree allows dynamic updates with higher maintenance overhead. Exact methods are valuable for small datasets, low dimensions, or strict accuracy needs. In large-scale high-dimensional settings, they are often replaced by approximate methods.
3.2 Approximate Nearest Neighbor Search
3.2.1 Hash-Based Approach
The core idea of the hash-based approach is to reduce search complexity by mapping high-dimensional data to lower-dimensional hash codes with carefully designed hash functions while preserving similarity between data points. As shown in Figure 3, each high-dimensional vector is transformed into a low-dimensional hash code. Similar points are mapped to the same or neighboring codes, so the search only needs to examine a small subset of codes, greatly improving efficiency. Based on this idea, four representative methods will be introduced: locality-sensitive hashing, spectral hashing, Spherical Hashing, and deep hashing.
The idea is to reduce the memory footprint and the search time by comparing the binary codes instead of the original vectors [cai2019revisit].
@@ 3.2.1 Hash-Based Approach
Local-Sensitive Hashing [datar2004locality, jafari2021survey]. It is a technique for finding the approximate nearest neighbors of a given vector in a large collection of vectors. It works by using a hash function to transform the high-dimensional vectors into compact binary codes, and then using a hash table to store and retrieve the codes based on their similarity or distance. In LSH, hash functions are designed to preserve the locality of vectors. Unlike traditional hash functions, LSH increases the probability that similar items are mapped to the same code, thus increasing collisions among similar vectors.
A trace of algorithm description and implementation for locally sensitive hashing can be seen on the home page [andoni2023lsh].
The LSH algorithm works by using a family of hash functions that use random projections or other techniques which are locality sensitive, meaning that similar vectors are more likely to have the same or similar codes than dissimilar vectors [dikkala2021manifold], which satisfy the following property:
@@
$f(d(p,q))=1-\frac{d(p,q)}{\pi w},$ (10)
where $d(p,q)$ is the Euclidean distance between $p$ and $q$. The advantage of LSH is that it can reduce the memory footprint and the search time by comparing the binary codes instead of the original vectors and also adapt to dynamic data sets by inserting or deleting codes from the hash table without affecting the existing codes [bob2022locality]. The performance of LSH depends on several factors, such as the dimensionality of the data, the number of hash functions, the number of bits per code, and the desired accuracy and recall. There are also some challenges and extensions of LSH, such as dealing with noisy and outlier data, choosing a good hash function family, or using multiple hash tables to increase recall. It is improved by [andoni2008near, andoni_beyond_2013, andoni_optimal_2015].
Spectral Hashing [weiss2008spectral]. It is a technique for finding the approximate nearest neighbors of a given vector in a large collection of vectors. It works by using spectral graph theory to generate hash functions that minimize the quantization error and maximize the variance of the binary codes. Spectral hashing can perform well when the data points lie on a low-dimensional manifold embedded in a high-dimensional space.
@@
$\vskip-2.84526pt\min_{y_{1},\ldots,y_{n}}\sum_{i=1}^{n}Var\left(y_{i}\right)-\lambda I\left(y_{1},\ldots,y_{n}\right)$ (11)
@@
Spherical Hashing. Spherical hashing is a binary encoding technique based on hyperspheres, designed for efficient ANNS. Unlike traditional hyperplane-based methods, it partitions the data space using hyperspheres, which define tighter and more compact regions through their centers and radii. Each spherical hashing function, as described by Heo [heo2012spherical], is characterized by $(p_{k}\in\mathbb{R}^{D})$ and a distance threshold $t_{k}\in\mathbb{R}^{+}$, as detailed below:
$\vskip-2.84526pth_{k}(x)=\begin{cases}-1&\text{ when }d\left(p_{k},x\right)>t_{k}\\ +1&\text{ when }d\left(p_{k},x\right)\leq t_{k}\end{cases}$ (12)
@@
Deep Hashing [liu2016deep, luo2022survey]. It is a technique for finding the approximate nearest neighbors of a given vector in a large collection of vectors. It works by using a deep neural network to learn hash functions that transform high-dimensional vectors into compact binary codes, and then using a hash table to store and retrieve the codes based on their similarity or distance [luo2023survey]. The hash functions are designed to preserve the semantic information of the vectors, which means that similar vectors are more likely to have the same or similar codes than dissimilar vectors [noauthor_remote_nodate].
@@
$\vskip-2.84526pt\min_{W,B}\sum_{i=1}^{N}\left\|x_{i}-Wb_{i}\right\|_{2}^{2}+\lambda\left\|b_{i}-\operatorname{sgn}\left(b_{i}\right)\right\|_{2}^{2}$ (14)
@@
The advantage of deep hashing is that it can leverage the representation learning ability of neural networks to generate more discriminative and robust codes for complex data, such as images, texts, or audio. The performance of deep hashing depends on several factors, such as the architecture of the neural network, the loss function used to train the network, and the number of bits per code.
3.2.2 Tree-Based Approach
The main idea of the tree-based approach is to build hierarchical or recursively partitioned data structures, such as trees, to break high-dimensional datasets into smaller subsets. This method improves query efficiency by reducing the number of points that need to be searched. Along this line, three tree-based methods will be presented: approximate nearest neighbors oh yeah, best bin first, and k-means tree.
The idea is to reduce the search space by following the branches of the tree that are most likely to contain the nearest neighbors of the query point.
Approximate Nearest Neighbors Oh Yeah [bernhardsson2015annoy]. It is a technique that can perform fast and accurate similarity searches and retrieval of high-dimensional vectors. It works by building a forest of binary trees, where each tree splits the vector space into two regions based on a random hyperplane. Each vector is then assigned to a leaf node in each tree based on which side of the hyperplane it falls on. To query a vector, Annoy traverses each tree from the root to the leaf node that contains the vector and collects all the vectors in the same leaf nodes as candidates. Then, it computes the exact distance or similarity between the query vector and each candidate and returns the top $k$ nearest neighbors. The formula for finding the median hyperplane between two points $p$ and $q$ is: